# On the Linear Speedup of the Push-Pull Method for Decentralized Optimization over Digraphs

### v1

@@ B.4 Proof of Lemma 6
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq s_{B}^{2}\|{\mathbf{g}}^{(0)}\|+s_{B}^{2}\sum_{k=0}^{T}\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}$
By Cauchy-Schwarz Inequality and our assumption for bounded noise, we have
$\displaystyle\mathbb{E}[\|{\mathbf{g}}^{(k+1)}-{\mathbf{g}}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 6n\sigma^{2}+3\mathbb{E}[\|\nabla f({\mathbf{x}}^{(k+1)})-\nabla f({\mathbf{x}}^{(k)})\|_{F}^{2}]$
$\displaystyle\leq 6n\sigma^{2}+3L^{2}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]$
Besides, $\mathbb{E}[\|{\mathbf{g}}^{(0)}\|_{F}^{2}]\leq n\sigma^{2}+\|\nabla f({\mathbf{x}}^{(0)})\|_{F}^{2}\leq n\sigma^{2}+2nL\Delta$. Therefore, we have
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7)\sigma^{2}+2ns_{B}^{2}L\Delta+3s_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}].$ (29)
Since
$\displaystyle\quad\mathbb{E}[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|_{F}^{2}]=\mathbb{E}[\|(A-I){\mathbf{x}}^{(k)}-\alpha{\mathbf{y}}^{(k)}\|_{F}^{2}]$ (30)
$\displaystyle=\mathbb{E}[\|(A-I)\Delta_{x}^{(k)}-\alpha\Delta_{y}^{(k)}+n\alpha\pi_{B}\bar{g}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq 3\|A-I\|_{2}^{2}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+3\alpha^{2}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]+3n\alpha^{2}\sigma^{2}+3n^{2}\alpha^{2}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}],$
and $\|A-I\|_{2}^{2}\leq 2n$, we can apply (30) to substitute the last term in (29). This produces
$\displaystyle(1-9\alpha^{2}L^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T)\sigma^{2}$
$\displaystyle\quad+9ns_{B}^{2}L^{2}\sum_{k=0}^{T}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+9\alpha^{2}L^{2}n^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]+2ns_{B}^{2}L\Delta$
Using the conclusion of Lemma 8, this can be further transformed to
$\displaystyle\quad(1-9\alpha^{2}L^{2}s_{B}^{2}-18\alpha^{2}L^{2}ns_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$
$\displaystyle\leq ns_{B}^{2}(6T+7+9\alpha^{2}L^{2}T+18\alpha^{2}L^{2}ns_{A}^{2}T)\sigma^{2}+2ns_{B}^{2}L\Delta$
$\displaystyle\quad+(9\alpha^{2}L^{2}n^{2}s_{B}^{2}+18\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2})\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
When $\alpha\leq\frac{1}{6\sqrt{n}s_{A}s_{B}L}$, this can be simplified to
$\displaystyle\sum_{k=0}^{T}\|\Delta_{y}^{(k+1)}\|_{F}^{2}\leq 30ns_{B}^{2}(T+1)\sigma^{2}+8ns_{B}^{2}L\Delta+108\alpha^{2}L^{2}n^{3}s_{A}^{2}s_{B}^{2}\sum_{k=0}^{T}\mathbb{E}[\|\overline{\nabla f}^{(k)}\|_{F}^{2}]$
B.5 Proof of Theorem 1
Plug Lemma 5 into Lemma 4, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\|\nabla f(\hat{x}^{(k)})\|^{2}+\frac{\alpha c(1-4\alpha mcL)}{2}\sum_{i=0}^{m-1}\mathbb{E}_{k}[\|\overline{\nabla f}^{(k+i)}\|^{2}]$ (31)
$\displaystyle\leq f(\hat{x}^{(k)})-\mathbb{E}_{k}[f(\hat{x}^{(k+m)})]+\left(\frac{2\alpha^{2}c^{2}mL}{n}+4\alpha^{2}Lns_{B}^{2}\right)\sigma^{2}$
$\displaystyle\quad+\frac{s_{A}^{2}L}{m}\|\Delta_{x}^{(k)}\|_{F}^{2}+4\alpha^{2}Ls_{B}^{2}\|\Delta_{y}^{(k)}\|_{F}^{2}.$
When $\alpha\leq\frac{1}{8cmL}$, $\frac{\alpha c(1-4\alpha mcL)}{2}$ can be replaced with $\frac{c\alpha}{4}$. Sum up (31) for $k\in[m^{*}]$, take full expectation and divide both sides by $\frac{c\alpha T}{4}$ we obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (32)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{4\alpha Lns_{B}^{2}}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{4s_{A}^{2}L}{cm\alpha T}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]+\frac{16\alpha Ls_{B}^{2}}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{x}^{(k+1)}\|_{F}^{2}$, we can plug in Lemma 8 into (32), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (33)
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+\left(\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}\right)\sigma^{2}$
$\displaystyle\quad+\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]+\frac{16\alpha L(s_{A}^{4}+s_{B}^{2})}{cmT}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}].$
Similarly, since $\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq\sum_{k=0}^{T-1}\|\Delta_{y}^{(k+1)}\|_{F}^{2}$, we further plug in Lemma 9 into (33), and obtain that
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (34)
@@ B.5 Proof of Theorem 1
$\displaystyle\leq\frac{4\Delta}{c\alpha T}+C_{\sigma}\sigma^{2}+\frac{128ns_{B}^{2}\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}+C_{f}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}].$
@@
$\displaystyle C_{\sigma}$ $\displaystyle=\frac{2c\alpha L}{n}+\frac{8\alpha Ln(s_{A}^{4}+s_{B}^{2})}{cm}+\frac{480\alpha Lns_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cm},$
$\displaystyle C_{f}$ $\displaystyle=\frac{8\alpha Ln^{2}s_{A}^{4}}{cmT}+\frac{2000\alpha^{3}L^{3}n^{3}s_{A}^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})}{cmT}.$
When $m\geq 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$, we have $C_{\sigma}\leq\frac{30c\alpha L}{n}$ and $C_{f}\leq\frac{2c\alpha L}{5T}+\frac{100c\alpha^{3}L^{3}n^{3}s_{A}^{2}}{T}$. When we further set $\alpha\leq\min\{\frac{1}{8cL},\frac{1}{4ns_{A}L}\}$, we have $C_{f}\leq\frac{1}{T}$. Under these conditions, (34) can be simplified to
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{4\Delta}{c\alpha T}+\frac{30c\alpha L}{n}\sigma^{2}+\frac{64n\alpha L^{2}(s_{A}^{4}+s_{B}^{2})\Delta}{cmT}.$ (35)
Finally, to satisfy all the constraints on $m$ and $\alpha$, we can set $m=\lceil 20n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})\rceil$, and
$\alpha=\min\left\{\frac{c}{21n^{2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})L},\frac{1}{8cL},\frac{\sqrt{n\Delta}\sigma}{3c\sqrt{LT}}\right\}.$
Plugging this into (35) gives that
$\displaystyle\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{20\sigma\sqrt{L\Delta}}{\sqrt{nT}}+C_{\Delta}\frac{L\Delta}{T},$
where
$C_{\Delta}=32+32ncs_{A}+84n^{2}c^{-2}s_{B}^{2}(s_{A}^{4}+s_{B}^{2})$
is a constant decided by $A$ and $B$.

### v2

@@ B.2 Proof of Lemma 5
$\displaystyle\quad\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k+m)}\|_{F}^{2}]\leq\alpha^{2}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\mathbb{E}_{k}[\|\Delta_{4}^{(k)}\|_{F}^{2}]]$
$\displaystyle\leq 15\alpha^{2}ns_{A}s_{B}s_{A^{m}}^{2}K\sigma^{2}+3\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
$\displaystyle\quad+6\alpha^{2}L^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]+3\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}\sum_{k\in[m^{*}]}\mathbb{E}\|\Delta_{y}^{(k)}\|_{F}^{2}$
Note that we start from consensual $x$, so $\Delta_{x}^{(0)}=0$. Plug in (37) and (36) we have
$\displaystyle\quad(1-480\alpha^{2}mL^{2}s_{A}s_{B}s_{A^{m}}^{2}n-960\alpha^{2}L^{2}ms_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2}n)\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]$
$\displaystyle\leq(15+126c^{2}\alpha^{2}L^{2}m^{2}n^{-2}+48s_{B}s_{B^{m}}^{2})\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}nK\sigma^{2}$
$\displaystyle\quad+(3+(126+480s_{B}s_{B^{m}}^{2})c^{2}\alpha^{2}L^{2}m^{2})\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]$
$\displaystyle\quad+(36+2880\alpha^{2}L^{2})s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}nL\Delta$
Using $\alpha\leq\min\{\frac{1}{10\sqrt{s_{B}s_{B^{m}}^{2}}cmL},\frac{1}{6L}\}$ and $m\geq 20c^{-2}ns_{A}s_{B}s_{A^{m}}^{2}$, the coefficients can be simplified to
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{x}^{(k)}\|_{F}^{2}]\leq 64\alpha^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2}nK\sigma^{2}$ (39)
$\displaystyle\quad+7\alpha^{2}n^{2}s_{A}s_{A^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+416\alpha^{2}s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}nL\Delta$
Taking (39) back to (36), we obtain that
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]\leq(1+1280\alpha^{2}L^{2}mn^{2}s_{A}s_{B}^{2}s_{B^{m}}^{2}s_{A^{m}}^{2})16s_{B}s_{B^{m}}^{2}Kn\sigma^{2}$
$\displaystyle\quad+(12+133120\alpha^{2}L^{2}s_{A}s_{B}^{2}s_{B^{m}}s_{A^{m}}^{2})s_{B^{m}}nL\Delta$
$\displaystyle\quad+160(c^{2}m^{2}s_{B}s_{B^{m}}^{2}+140\alpha^{2}L^{2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}s_{B^{m}}^{2})\alpha^{2}L^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq 10c^{-2}n^{2}s_{A}s_{A^{m}}^{2}s_{B}$, we have
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$ $\displaystyle\leq 32s_{B}s_{B^{m}}^{2}Kn\sigma^{2}+26s_{B^{m}}nL\Delta$ (40)
$\displaystyle\quad+170c^{2}m^{2}\alpha^{2}L^{2}s_{B}s_{B^{m}}^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
Plug (39) into (37), we obtain that
$\displaystyle\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{2}^{(k)}]\leq(21c^{2}m^{2}n^{-2}+3200mns_{A}s_{B}^{2}s_{A^{m}}^{2}s_{B^{m}}^{2})\alpha^{2}nK\sigma^{2}$ (41)
$\displaystyle\quad+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+(21c^{2}m+350n^{3}s_{A}s_{A^{m}}^{2})\alpha^{2}m\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}].$
$\displaystyle\leq 181c^{2}m^{2}n^{-1}\alpha^{2}K\sigma^{2}+21280s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}mn^{2}L\Delta+38\alpha^{2}c^{2}m^{2}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}],$
where the second inequality uses $m\geq 20c^{-2}n^{3}s_{A}s_{A^{m}}^{2}s_{B}^{2}s_{B^{m}}^{2}$. Now we are ready to estimate the summation of error terms and prove Lemma 5. Remember that $\Delta_{1}^{(k)}$ and $\Delta_{2}^{(k)}$ share the same upper bound, we utilize (41), (40) and obtain that
$\displaystyle\frac{c\alpha L^{2}}{2}\quad\sum_{k\in[m^{*}]}\mathbb{E}[\Delta_{1}^{(k)}+\Delta_{2}^{(k)}]+\frac{3\alpha s_{B}^{2}}{cm}\sum_{k\in[m^{*}]}\mathbb{E}[\|\Delta_{y}^{(k)}\|_{F}^{2}]$
@@
$\displaystyle\leq\frac{\alpha cm}{4}C_{\sigma,1}K\sigma^{2}+\frac{c\alpha}{4}C_{f,1}\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|_{F}^{2}]+C_{\Delta,1}\frac{c\alpha L\Delta}{4m},$
@@
$\displaystyle C_{\sigma,1}$ $\displaystyle=724c^{2}mn^{-1}\alpha^{2}L^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$
$\displaystyle C_{f,1}$ $\displaystyle=152c^{2}\alpha^{2}L^{2}m^{2}+2040m\alpha^{2}L^{2}s_{B}^{3}s_{B^{m}}^{2}$
$\displaystyle C_{\Delta,1}$ $\displaystyle=85120s_{A}s_{B}s_{A^{m}}^{2}s_{B^{m}}\alpha^{2}L^{2}m^{2}n^{2}+\frac{312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
When $\alpha\leq\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L}$ and $m\geq c^{-2}30s_{B}^{1.5}$, we have
$C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}},\quad C_{f,1}\leq 1,\quad C_{\Delta,1}\leq\frac{10s_{A}s_{A^{m}}^{2}n^{2}+312s_{B}^{2}s_{B^{m}}n}{c^{2}}.$
This finishes the proof.
B.3 Proof of Theorem 1
Sum up Lemma 4 for $k\in[m^{*}]$, then plug in Lemma 5, we obtain that
$\displaystyle\quad\frac{\alpha cm}{4}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]+\frac{\alpha c}{4}(1-C_{f,1})\sum_{t=0}^{T-1}\mathbb{E}[\|\overline{\nabla f}^{(t)}\|^{2}]$ (42)
$\displaystyle\leq f(x^{(0)})-f(\hat{x}^{T})+\frac{\alpha cm}{4}\left(\frac{8\alpha cL}{n}+\frac{8\alpha Lns_{B}^{2}}{cm}+C_{\sigma,1}\right)K\sigma^{2}+C_{\Delta,1}\frac{c\alpha L\Delta}{4m}.$
Divide both sides by $\frac{\alpha cmK}{4}$ and utilize $C_{f,1}\leq 1$, $f(x^{(0)})-f(\hat{x}^{T})\leq\Delta$, $m\geq c^{-2}n^{2}s_{B}^{2}$ and $C_{\sigma,1}\leq\frac{73c\alpha L}{n}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}$, $T=mK$, we have
$\displaystyle\quad\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]\leq\frac{\Delta}{c\alpha T}+\frac{100c\alpha L}{n}\sigma^{2}+\frac{384s_{B}^{3}s_{B^{m}}^{2}}{m^{2}c^{2}}\sigma^{2}+C_{\Delta,1}\frac{L\Delta}{mT}.$ (43)
Inequality (43) holds when $\alpha\leq\min\{\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}},\frac{1}{6L}\}$ and $m$ is larger than some matrix-related constants.
To obtain a standard linear speedup result, we set
$\displaystyle m$ $\displaystyle=\lceil\max\{4c^{-1}s_{B}^{1.5}s_{B^{m}}\left(\frac{nT\sigma^{2}}{L\Delta}\right)^{0.25},C_{m}\}\rceil,$
$\displaystyle\alpha$ $\displaystyle=\min\left\{\frac{\sqrt{n\Delta}\sigma}{10c\sqrt{LT}},\frac{1}{10cm\sqrt{s_{B}s_{B^{m}}^{2}}L},\frac{1}{6L}\right\},$
giving that
$\displaystyle\quad\min_{t\in\{0,1,\ldots,T-1\}}\mathbb{E}[\|\nabla f(\hat{x}^{(t)}\|^{2}]\leq\frac{1}{K}\sum_{k\in[m^{*}]}\mathbb{E}[\|\nabla f(\hat{x}^{(k)})\|^{2}]$
$\displaystyle\leq\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+\frac{L\Delta(m^{-1}C_{\Delta,1}+10m\sqrt{s_{B}s_{B^{m}}^{2}}+c^{-1})}{T}$
$\displaystyle=\frac{44\sigma\sqrt{L\Delta}}{\sqrt{nT}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)^{\frac{3}{4}}+{\mathcal{O}}\left(\frac{L\Delta}{T}\right)$